OpenStudy (anonymous):

‘Tom travels from a to b and at a speed of 90km per hour. It takes him time ‘t’. Alice travels from a to b at a speed of 70km per hour. It takes her time ‘t+ 1 hour’. How long did it take for Tom to complete his journey?’

OpenStudy (anonymous):

@asnaseer

OpenStudy (asnaseer):

Call the distance between a and b 'd'. This distance does not change. Now calculate 'd' based on Toms speed and time. Then calculate 'd' based on Alice's speed and time. Both values should be equal. Use this fact to solve for 't' and that should lead you to the answer.

OpenStudy (anonymous):

so 3.5

OpenStudy (asnaseer):

Please show your steps so that I can spot where you may have made a mistake

OpenStudy (anonymous):

90t=70(t+1) 20t=70 t=70/20=3.5 hours

OpenStudy (asnaseer):

Oh - ok - that IS correct - I /thought/ you had given the value for distance between a and b :)